OpenStudy (anonymous):

the sleigh have the weight 50N and slides downhill in a constant speed. a) name all the forces that is affecting the sleigh b) calculate the friction to magnitude and direction

OpenStudy (anonymous):



OpenStudy (anonymous):

a) = Ff, FN & g? But b) I need help with

OpenStudy (anonymous):

Friction is defined by the following equation:\[F_{f}=\mu N\]'mu' is the frictional coefficient and 'N' is the normal force. There is a problem though. We don't know 'mu'. So, we need another method to figure this out. We are told the sled slides at constant speed. That means there are no net forces acting on the sled. By this, we can say that the force due to gravity acting along the slope must be equal to the force of friction. We know the the component of gravity acting from geometry: So we can say that this is the magnitude of the frictional force acting up the slope:\[F_{f}=mg \sin(\theta)=W\sin(\theta)\]Sine of theta is actually given by the ratio of the lengths provided:\[\sin(\theta)=\frac{0.75}{1.5}=\frac{1}{2}\]\[F_{f}=W(\frac{1}{2})=50(\frac{1}{2})\] Hence,\[F_{f}=25N\]directed upwards, parallel to the slope

OpenStudy (anonymous):

thanks alot agaim @Pompeii00 :D